# Bias and Fairness in Large Language Models: A Survey

### v2

@@ 3.3 Embedding-Based Metrics
In this section, we discuss bias evaluation metrics that leverage embeddings. Embedding-based metrics typically compute distances in the vector space between neutral words, such as professions, and identity-related words, such as gender pronouns. We present one relevant method for static word embeddings, and focus otherwise on sentence-level contextualized embeddings used in LLMs. We illustrate an example of these metrics in Figure 3.
Figure 3: Example Embedding-Based Metrics (§ 3.3). Sentence-level encoders produce sentence embeddings that can be assessed for bias. Embedding-based metrics use cosine similarity to compare words like "doctor" to social group terms like "man". Unbiased embeddings should have similar cosine similarity to opposing social group terms.
@@ 3.3.1 Word Embedding Metrics
Bias metrics for word embeddings were first proposed for static word embeddings, but their basic formulation of computing cosine distances between neutral and gendered words has been generalized to contextualized embeddings and broader dimensions of bias. Static embedding techniques may be adapted to contextualized embeddings by taking the last subword token representation of a word before pooling to a sentence embedding. Though several static word embedding bias metrics have been proposed, we focus only on Word Embedding Association Test (WEAT) (Caliskan et al., 2017) here, given its relevance to similar methods for contextualized sentence embeddings. WEAT measures associations between social group concepts (e.g., masculine and feminine words) and neutral attributes (e.g., family and occupation words), emulating the Implicit Association Test (Greenwald et al., 1998). For protected attributes $A_{1}$, $A_{2}$ and neutral attributes $W_{1}$, $W_{2}$, stereotypical associations are measured by a test statistic:
@@
with a larger effect size indicating stronger bias. WEAT* (Dev et al., 2021) presents an alternative, where $W_{1}$ and $W_{2}$ are instead definitionally masculine and feminine words (e.g., "gentleman", "matriarch") to capture stronger masculine and feminine associations.
@@ 3.3.2 Sentence Embedding Metrics
Several of these methods follow WEAT’s formulation. To adapt WEAT to contextualized embeddings, Sentence Encoder Association Test (SEAT) (May et al., 2019) generate embeddings of semantically bleached template-based sentences (e.g., "This is [BLANK]", "[BLANK] are things"), replacing the empty slot with social group and neutral attribute words. The same formulation in Equation 5 applies, using the [CLS] token as the embeddings. SEAT can be extended to measure more specific dimensions of bias with unbleached templates, such as, "The engineer is [BLANK]." Tan & Celis (2019) similarly extend WEAT to contextualized embeddings by extracting contextual word embeddings before they are pooled to form a sentence embedding.
Contextualized Embedding Association Test (CEAT) (Guo & Caliskan, 2021) uses an alternative approach to extend WEAT to contextualized embeddings. Instead of calculating WEAT’s effect size given by Equation 5 directly, it generates sentences with combinations of $A_{1}$, $A_{2}$, $W_{1}$, and $W_{2}$, randomly samples a subset of embeddings, and calculates a distribution of effect sizes. The magnitude of bias is calculated with a random-effects model, and is given by:
@@
Instead of using the sentence-level representation, Sentence Bias Score (Dolci et al., 2023) computes a normalized sum of word-level biases. The metric uses the encodings of an original sentence and its counterfactual with a binary protected attribute swapped. The cosine similarity between word embeddings and a gender direction $\bm{\mathrm{g}}$ in the embedding space produces word-level bias. Therefore, for semantic importance $\alpha_{s}$ of each word $s$ in a sentence $S$, the sentence bias is given by:
Sentence Bias($\textrm{Sentence Bias}(S)=\sum_{s\in S}|\cos(\bm{\mathrm{s}},\bm{\mathrm{g}})\cdot\alpha_{s}|$ (7)
@@ 3.3.3 Discussion and Limitations
Several works point out that biases in the embedding space have only weak or inconsistent relationships with biases in downstream tasks (Cabello et al., 2023; Cao et al., 2022; Goldfarb-Tarrant et al., 2021; Orgad & Belinkov, 2022; Orgad et al., 2022; Steed et al., 2022). In fact, Goldfarb-Tarrant et al. (2021) find no reliable correlation at all, and Cabello et al. (2023) illustrate that associations between the representations of protected attribute and other words can be independent of downstream performance disparities, if certain assumptions of social groups’ language use are violated. These works demonstrate that bias in representations and bias in downstream applications should not be conflated, which may limit the value of embedding-based metrics. Delobelle et al. (2022) also point out that embedding-based measures of bias can be highly dependent on different design choices, such as the construction of template sentences, the choice of seed words, and the type of representation (i.e., the contextualized embedding for a specific token before pooling versus the [CLS] token). In fact, Delobelle et al. (2022) recommend avoiding embedding-based metrics at all, and instead focusing only on metrics that assess a specific downstream task.
Furthermore, Gonen & Goldberg (2019) critically show that debiasing techniques may merely represent bias in new ways in the embedding space. This finding may also call the validity of embedding-based metrics into question. Particularly, whether embedding-based metrics, with their reliance on cosine distance, sufficiently capture only superficial levels of bias, or whether they can also identify more subtle forms of bias, is a topic for future research.
@@ 3.4 Probability-Based Metrics
Figure 4: Example Probability-Based Metrics (§ 3.4). We illustrate two classes of probability-based metrics: masked token metrics and pseudo-log-likelihood metrics. Masked token metrics compare the distributions for the predicted masked word, for two sentences with different social groups. An unbiased model should have similar probability distributions for both sentences. Pseudo-log-likelihood metrics estimate if a stereotype or anti-stereotype sentence is more likely by approximating the conditional probability of the sentence given each word in the sentence. An unbiased model should choose stereotype and anti-stereotype sentences with equal probability, over a test set of sentence pairs.
@@ 3.4.1 Masked Token Methods
The probability of a token can be derived by masking a word in a sentence and asking a masked language model to fill in the blank. Discovery of Correlations (DisCo) (Webster et al., 2020), for instance, compares completions of template sentences. Each template (e.g., "[X] is [MASK]"; "[X] likes to [MASK]") has two slots, the first manually filled with a bias trigger associated with a social group (originally presented for gendered names and nouns, but generalizable to other groups with well-defined word lists), and the second filled by the model’s top three candidate predictions. The score is calculated by averaging the count of differing predictions between social groups across all templates. Log-Probability Bias Score (LPBS) (Kurita et al., 2019) uses a similar template-based approach as DisCo to measure bias in neutral attribute words (e.g., occupations), but normalizes a token’s predicted probability $p_{a}$ (based on a template "[MASK] is a [NEUTRAL ATTRIBUTE]") with the model’s prior probability $p_{prior}$ (based on a template "[MASK] is a [MASK]"). Normalization corrects for the model’s prior favoring of one social group over another and thus only measures bias attributable to the [NEUTRAL ATTRIBUTE] token. Bias is measured by the difference between normalized probability scores for two binary and opposing social group words.
@@
Categorical Bias Score (Ahn & Oh, 2021) adapts Kurita et al. (2019)’s normalized log probabilities to non-binary targets. This metric measures the variance of predicted tokens for fill-in-the-blank template prompts over corresponding protected attribute words $a$ for different social groups:
@@ 3.4.2 Pseudo-Log-Likelihood Methods
Several techniques leverage pseudo-log-likelihood (PLL) (Salazar et al., 2020; Wang & Cho, 2019) to score the probability of generating a token given other words in the sentence. For a sentence $S$, PLL is given by:
@@
Context Association Test (CAT) (Nadeem et al., 2021), introduced with the StereoSet dataset, also compares sentences. Similar to pseudo-log-likelihood, each sentence is paired with a stereotype, anti-stereotype, and meaningless option, which are either fill-in-the-blank tokens or continuation sentences. In contrast to pseudo-log-likelihood, CAT considers $P(M|U,\theta)$, rather than $P(U|M,\theta)$. This can be framed as:
@@
Idealized CAT (iCAT) Score can be calculated from the same stereotype, anti-stereotype, and meaningless sentence options. Nadeem et al. (2021) define an idealized language model to have a language modeling score ($lms$) equal to 100 (i.e., it always chooses a meaningful option), and a stereotype score ($ss$) of 50 (i.e., it chooses an equal number of stereotype and anti-stereotype options).
@@
All Unmasked Likelihood (AUL) (Kaneko & Bollegala, 2022) extends the CrowS-Pair Score and CAT to consider multiple correct candidate predictions. While pseudo-log-likelihood and CAT consider a single correct answer for a masked test example, AUL provides an unmasked sentence to the model and predicts all tokens in the sentence. The unmasked input provides the model with all information to predict a token, which can improve the prediction accuracy of the model, and avoids selection bias in the choice of which words to mask.
@@
Kaneko & Bollegala (2022) also provides a variation dubbed AUL with Attention Weights (AULA) that considers attention weights to account for different token importances. With $\alpha_{i}$ as the attention associated with $s_{i}$, AULA is given by:
@@ 3.4.3 Discussion and Limitations
Similar to the shortcomings of embedding-based metrics, Delobelle et al. (2022) and Kaneko et al. (2022) point out that probability-based metrics may be only weakly correlated with biases that appear in downstream tasks, and caution that these metrics are not sufficient checks for bias prior to deployment. Thus, probability-based metrics should be paired with additional metrics that more directly assess a downstream task.
Each class of probability-based metrics also carries some risks. Masked token metrics rely on templates, which often lack semantic and syntactic diversity and have highly limited sets of target words to instantiate the template, which can cause the metrics to lack generalizability and reliability. Blodgett et al. (2021) highlight shortcomings of pseudo-log-likelihood metrics that compare stereotype and anti-stereotype sentences. The notion that stereotype and anti-stereotype sentences should be selected at equal rates (using Equation 16) is not obvious as an indicator of fairness, and may depend heavily on the conceptualization of what stereotypes and anti-stereotypes entail in the evaluation dataset. Furthermore, merely selecting between two sentences may not fully capture the tendency of a model to produce stereotypical outputs, and can misrepresent the model’s behavior by ranking sentences instead of more carefully examining the magnitude of likelihoods directly.
@@ 3.5 Generated Text-Based Metrics
Figure 5: Example Generated Text-Based Metrics (§ 3.5). Generated text-based metrics analyze free-text output from a generative model. Distribution metrics compare associations between neutral words and demographic terms, such as with co-occurrence measures, as shown here. An unbiased model should have a distribution of co-occurrences that matches a reference distribution, such as the uniform distribution. Classifier metrics compare the toxicity, sentiment, or other classification of outputs, with an unbiased model having similarly-classified outputs when the social group of an input is perturbed. Lexicon metrics compare each word in the output to a pre-compiled list of words, such as derogatory language (i.e., "@&!", "#\$!") in this example, to generate a bias score. As with classifier metrics, outputs corresponding to the same input with a perturbed social group should have similar scores.
@@ 3.5.1 Distribution Metrics
This metric may be overly stringent, however. Other metrics instead look at the distribution of terms that appear nearby social group terms. One common measure is the Co-Occurrence Bias Score (Bordia & Bowman, 2019), which measures the co-occurrence of tokens with gendered words in a corpus of generated text. For a token $w$ and two sets of attribute words $A_{i}$ and $A_{j}$, the bias score for each word is given by:
@@
with a score of zero for words that co-occur equally with feminine and masculine gendered words. In a similar vein, Demographic Representation (DR) (Bommasani et al., 2023) compares the frequency of mentions of social groups to the original data distribution. Let $C(x,Y)$ be the count of how many times word $x$ appears in the sequence $Y$. For each group $G_{i}\in\mathbb{G}$ with associated protected attribute words $A_{i}$, the count $\textrm{DR}(G_{i})$ is
@@
The vector of counts $\textrm{DR}=[\textrm{DR}(G_{1}),\ldots,\textrm{DR}(G_{m})]$ normalized to a probability distribution can then be compared to a reference probability distribution (e.g., uniform distribution) with metrics like total variation distance, KL divergence, Wasserstein distance, or others. Stereotypical Associations (ST) (Bommasani et al., 2023) measures bias associated with specific terms, defined as:

### v3

@@ 3.3 Embedding-Based Metrics
In this section, we discuss bias evaluation metrics that leverage embeddings. Embedding-based metrics typically compute distances in the vector space between neutral words, such as professions, and identity-related words, such as gender pronouns. We present one relevant method for static word embeddings, and focus otherwise on sentence-level contextualized embeddings used in LLMs. We illustrate an example in Figure 3.
Figure 3: Example Embedding-Based Metrics (§ 3.3). Sentence-level encoders produce sentence embeddings that can be assessed for bias. Embedding-based metrics use cosine similarity to compare words like "doctor" to social group terms like "man." Unbiased embeddings should have similar cosine similarity to opposing social group terms.
@@ 3.3.1 Word Embedding Metrics
Bias metrics for word embeddings were first proposed for static word embeddings, but their basic formulation of computing cosine distances between neutral and gendered words has been generalized to contextualized embeddings and broader dimensions of bias. Static embedding techniques may be adapted to contextualized embeddings by taking the last subword token representation of a word before pooling to a sentence embedding. Though several static word embedding bias metrics have been proposed, we focus only on Word Embedding Association Test (WEAT) (Caliskan, Bryson, and Narayanan, 2017) here, given its relevance to similar methods for contextualized sentence embeddings. WEAT measures associations between social group concepts (e.g., masculine and feminine words) and neutral attributes (e.g., family and occupation words), emulating the Implicit Association Test (Greenwald, McGhee, and Schwartz, 1998). For protected attributes $A_{1}$, $A_{2}$ and neutral attributes $W_{1}$, $W_{2}$, stereotypical associations are measured by a test statistic:
@@
with a larger effect size indicating stronger bias. WEAT* (Dev et al., 2021) presents an alternative, where $W_{1}$ and $W_{2}$ are instead definitionally masculine and feminine words (e.g., "gentleman," "matriarch") to capture stronger masculine and feminine associations.
@@ 3.3.2 Sentence Embedding Metrics
Several of these methods follow WEAT’s formulation. To adapt WEAT to contextualized embeddings, Sentence Encoder Association Test (SEAT) (May et al., 2019) generates embeddings of semantically bleached template-based sentences (e.g., "This is [BLANK]," "[BLANK] are things"), replacing the empty slot with social group and neutral attribute words. The same formulation in Equation 5 applies, using the [CLS] token as the embeddings. SEAT can be extended to measure more specific dimensions of bias with unbleached templates, such as, "The engineer is [BLANK]." Tan and Celis (2019) similarly extend WEAT to contextualized embeddings by extracting contextual word embeddings before they are pooled to form a sentence embedding.
Contextualized Embedding Association Test (CEAT) (Guo and Caliskan, 2021) uses an alternative approach to extend WEAT to contextualized embeddings. Instead of calculating WEAT’s effect size given by Equation 5 directly, it generates sentences with combinations of $A_{1}$, $A_{2}$, $W_{1}$, and $W_{2}$, randomly samples a subset of embeddings, and calculates a distribution of effect sizes. The magnitude of bias is calculated with a random-effects model, and is given by:
@@
Instead of using the sentence-level representation, Sentence Bias Score (Dolci, Azzalini, and Tanelli, 2023) computes a normalized sum of word-level biases. Given a sentence $S$ and a list of gendered words $A$, the metric computes the cosine similarity between the embedding of each word $s$ in the sentence $S$ and a gender direction $\bm{\mathrm{v}}_{\textrm{gender}}$ in the embedding space. The gender direction is identified by the difference between the embeddings of feminine and masculine gendered words, reduced to a single dimension with principal component analysis (PCA). The sentence importance weighs each word-level bias by a semantic importance score $\alpha_{s}$, given by the number of times the sentence encoder’s max-pooling operation selects the representation at $s$’s position $t$.
Sentence Bias($\textrm{Sentence Bias}(S)=\sum_{s\in S,s\notin A}|\cos(\bm{\mathrm{s}},\bm{\mathrm{v}}_{\textrm{gender}})\cdot\alpha_{s}|$ (7)
@@ 3.3.3 Discussion and Limitations
Several works point out that biases in the embedding space have only weak or inconsistent relationships with biases in downstream tasks (Cabello, Jørgensen, and Søgaard, 2023; Cao et al., 2022a; Goldfarb-Tarrant et al., 2021; Orgad and Belinkov, 2022; Orgad, Goldfarb-Tarrant, and Belinkov, 2022; Steed et al., 2022). In fact, Goldfarb-Tarrant et al. (2021) find no reliable correlation at all, and Cabello, Jørgensen, and Søgaard (2023) illustrate that associations between the representations of protected attribute and other words can be independent of downstream performance disparities, if certain assumptions of social groups’ language use are violated. These works demonstrate that bias in representations and bias in downstream applications should not be conflated, which may limit the value of embedding-based metrics. Delobelle et al. (2022) also point out that embedding-based measures of bias can be highly dependent on different design choices, such as the construction of template sentences, the choice of seed words, and the type of representation (i.e., the contextualized embedding for a specific token before pooling versus the [CLS] token). In fact, Delobelle et al. (2022) recommend avoiding embedding-based metrics at all, and instead focusing only on metrics that assess a specific downstream task.
Furthermore, Gonen and Goldberg (2019) critically show that debiasing techniques may merely represent bias in new ways in the embedding space. This finding may also call the validity of embedding-based metrics into question. Particularly, whether embedding-based metrics, with their reliance on cosine distance, sufficiently capture only superficial levels of bias, or whether they can also identify more subtle forms of bias, is a topic for future research.
@@ 3.4 Probability-Based Metrics
Figure 4: Example Probability-Based Metrics (§ 3.4). We illustrate two classes of probability-based metrics: masked token metrics and pseudo-log-likelihood metrics. Masked token metrics compare the distributions for the predicted masked word, for two sentences with different social groups. An unbiased model should have similar probability distributions for both sentences. Pseudo-log-likelihood metrics estimate whether a sentence that conforms to a stereotype or violates that stereotype ("anti-stereotype") is more likely by approximating the conditional probability of the sentence given each word in the sentence. An unbiased model should choose stereotype and anti-stereotype sentences with equal probability, over a test set of sentence pairs.
@@ 3.4.1 Masked Token Methods
The probability of a token can be derived by masking a word in a sentence and asking a masked language model to fill in the blank. Discovery of Correlations (DisCo) (Webster et al., 2020), for instance, compares the completion of template sentences. Each template (e.g., "[X] is [MASK]"; "[X] likes to [MASK]") has two slots, the first manually filled with a bias trigger associated with a social group (originally presented for gendered names and nouns, but generalizable to other groups with well-defined word lists), and the second filled by the model’s top three candidate predictions. The score is calculated by averaging the count of differing predictions between social groups across all templates. Log-Probability Bias Score (LPBS) (Kurita et al., 2019) uses a similar template-based approach as DisCo to measure bias in neutral attribute words (e.g., occupations), but normalizes a token’s predicted probability $p_{a}$ (based on a template "[MASK] is a [NEUTRAL ATTRIBUTE]") with the model’s prior probability $p_{prior}$ (based on a template "[MASK] is a [MASK]"). Normalization corrects for the model’s prior favoring of one social group over another and thus only measures bias attributable to the [NEUTRAL ATTRIBUTE] token. Bias is measured by the difference between normalized probability scores for two binary and opposing social group words.
@@
Categorical Bias Score (Ahn and Oh, 2021) adapts Kurita et al. (2019)’s normalized log probabilities to non-binary targets. This metric measures the variance of predicted tokens for fill-in-the-blank template prompts over corresponding protected attribute words $a$ for different social groups:
@@ 3.4.2 Pseudo-Log-Likelihood Methods
Several techniques leverage pseudo-log-likelihood (PLL) (Salazar et al., 2020; Wang and Cho, 2019) to score the probability of generating a token given other words in the sentence. For a sentence $S$, PLL is given by:
@@
Context Association Test (CAT) (Nadeem, Bethke, and Reddy, 2021), introduced with the StereoSet dataset, also compares sentences. Similar to pseudo-log-likelihood, each sentence is paired with a stereotype, "anti-stereotype," and meaningless option, which are either fill-in-the-blank tokens or continuation sentences. The stereotype sentence illustrates a stereotype about a social group, while the anti-stereotype sentence replaces the social group with an instantiation that violates the given stereotype; thus, anti-stereotype sentences do not necessarily reflect pertinent harms. In contrast to pseudo-log-likelihood, CAT considers $P(M|U,\theta)$, rather than $P(U|M,\theta)$. This can be framed as:
@@
Idealized CAT (iCAT) Score can be calculated from the same stereotype, anti-stereotype, and meaningless sentence options. Given a language modeling score ($lms$) that calculates the percentage of instances that the model prefers a meaningful sentence option over a meaningless one, as well as a stereotype score ($ss$) that calculates the percentage of instances that the model prefers a stereotype option over an anti-stereotype one, Nadeem, Bethke, and Reddy (2021) define an idealized language model to have a language modeling score equal to 100 (i.e., it always chooses a meaningful option) and a stereotype score of 50 (i.e., it chooses an equal number of stereotype and anti-stereotype options).
@@
All Unmasked Likelihood (AUL) (Kaneko and Bollegala, 2022) extends the CrowS-Pair Score and CAT to consider multiple correct candidate predictions. While pseudo-log-likelihood and CAT consider a single correct answer for a masked test example, AUL provides an unmasked sentence to the model and predicts all tokens in the sentence. The unmasked input provides the model with all information to predict a token, which can improve the prediction accuracy of the model, and avoids selection bias in the choice of which words to mask.
@@
Kaneko and Bollegala (2022) also provides a variation dubbed AUL with Attention Weights (AULA) that considers attention weights to account for different token importances. With $\alpha_{i}$ as the attention associated with $s_{i}$, AULA is given by:
@@ 3.4.3 Discussion and Limitations
Similar to the shortcomings of embedding-based metrics, Delobelle et al. (2022) and Kaneko, Bollegala, and Okazaki (2022) point out that probability-based metrics may be only weakly correlated with biases that appear in downstream tasks, and caution that these metrics are not sufficient checks for bias prior to deployment. Thus, probability-based metrics should be paired with additional metrics that more directly assess a downstream task.
Each class of probability-based metrics also carries some risks. Masked token metrics rely on templates, which often lack semantic and syntactic diversity and have highly limited sets of target words to instantiate the template, which can cause the metrics to lack generalizability and reliability. Blodgett et al. (2021) highlight shortcomings of pseudo-log-likelihood metrics that compare stereotype and anti-stereotype sentences. The notion that stereotype and anti-stereotype sentences, which, by construction, do not reflect real-world power dynamics, should be selected at equal rates (using Equation 16) is not obvious as an indicator of fairness, and may depend heavily on the conceptualization of what stereotypes and anti-stereotypes entail in the evaluation dataset (see further discussion in Section 4.1.3). Furthermore, merely selecting between two sentences may not fully capture the tendency of a model to produce stereotypical outputs, and can misrepresent the model’s behavior by ranking sentences instead of more carefully examining the magnitude of likelihoods directly.
@@ 3.5 Generated Text-Based Metrics
Figure 5: Example Generated Text-Based Metrics (§ 3.5). Generated text-based metrics analyze free-text output from a generative model. Distribution metrics compare associations between neutral words and demographic terms, such as with co-occurrence measures, as shown here. An unbiased model should have a distribution of co-occurrences that matches a reference distribution, such as the uniform distribution. Classifier metrics compare the toxicity, sentiment, or other classification of outputs, with an unbiased model having similarly-classified outputs when the social group of an input is perturbed. Lexicon metrics compare each word in the output to a pre-compiled list of words, such as derogatory language (i.e., "@&!," "#\$!") in this example, to generate a bias score. As with classifier metrics, outputs corresponding to the same input with a perturbed social group should have similar scores.
@@ 3.5.1 Distribution Metrics
This metric may be overly stringent, however. Other metrics instead look at the distribution of terms that appear nearby social group terms. One common measure is the Co-Occurrence Bias Score (Bordia and Bowman, 2019), which measures the co-occurrence of tokens with gendered words in a corpus of generated text. For a token $w$ and two sets of attribute words $A_{i}$ and $A_{j}$, the bias score for each word is given by:
@@
with a score of zero for words that co-occur equally with feminine and masculine gendered words. In a similar vein, Demographic Representation (DR) (Liang et al., 2022) compares the frequency of mentions of social groups to the original data distribution. Let $C(x,Y)$ be the count of how many times word $x$ appears in the sequence $Y$. For each group $G_{i}\in\mathbb{G}$ with associated protected attribute words $A_{i}$, the count $\textrm{DR}(G_{i})$ is
@@
The vector of counts $\textrm{DR}=[\textrm{DR}(G_{1}),\ldots,\textrm{DR}(G_{m})]$ normalized to a probability distribution can then be compared to a reference probability distribution (e.g., uniform distribution) with metrics like total variation distance, KL divergence, Wasserstein distance, or others. Stereotypical Associations (ST) (Liang et al., 2022) measures bias associated with specific terms, defined as:
@@
To measure linguistic differences between unmarked default, Cheng, Durmus, and Jurafsky (2023) presents Marked Persons, leveraging the concept of markedness to measure stereotypes, and comparing the marked language of marginalized groups (e.g., "Black woman") to the unmarked language of dominant, default ones (e.g., "white," "man"). After prompting a model to write a persona of a specified identity, this technique identifies words that statistically distinguish a marked group from an unmarked one.